Problem 2.12 Use Gauss's law find the electric field inside a uniformly charged sphere (charge density p). Compare your answer to Prob. 2.8.


1
Expert's answer
2021-09-26T18:48:29-0400

To find electric field inside a charged solid sphere using Gauss's law.

Consider a charged sphere of radius R and uniform charge density ρ, q is the charge on sphere where "ρ=\frac{q}{V}=\frac{q}{\frac{4}{3} \pi R^3}"

Consider a point P at a distance r inside the sphere.

Now draw a Gaussian surface with radius r<R



Using Gauss’s theorem

"Φ = \oint \vec{E}\cdot \vec{ds} = \frac{q_{enclosed}}{ε_0} \\

q_{enclosed} = ρV_{enclosed} = (\frac{4}{3} \pi R^3)(\frac{4}{3} \pi r^3) \\

q_{enclosed} = q(\frac{r^3}{R^3}) \\

\oint \vec{E}\cdot \vec{ds} = \frac{qr^3}{ε_0R^3} \\

E4 \pi r^2 = \frac{qr^3}{ε_0R^3} \\

E= \frac{1}{4 \pi ε_0} \frac{qr}{R^3}"

In terms of charge density

"E= \frac{1}{4 \pi ε_0} \frac{(ρ \frac{4}{3} \pi R^3)r}{R^3} \\

E = \frac{ρr}{3ε_0 }"